40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Fri Aug 17 14:56:59 2018, aaron, Update, Cryo Q, Finding Modes InAirGeNSTF.png
    Reply  Mon Aug 20 22:03:26 2018, Mandy , Update, Cryo Q, Finding Modes 
       Reply  Mon Aug 20 23:40:29 2018, aaron, Update, Cryo Q, Finding Modes 180820_TFscan.png180820_kHzModeTF.png
          Reply  Tue Aug 21 11:29:30 2018, aaron, Update, Cryo Q, Finding Modes 
             Reply  Wed Aug 22 10:07:03 2018, aaron, Update, Cryo Q, Finding Modes 
                Reply  Wed Aug 22 13:17:19 2018, aaron, Computing, Cryo Q, timing error, rebuilding model 
                   Reply  Wed Aug 22 15:48:35 2018, aaron, Computing, Cryo Q, timing error, rebuilding model phase_rotator.png1kHzrindown.png
                      Reply  Thu Aug 23 11:14:19 2018, aaron, Computing, Cryo Q, rebuilding model, moderinger MRChain_woExcitation.pngMRChain_wExcitation.png
                   Reply  Wed Aug 22 21:55:02 2018, johannes, Computing, stuff happens, problem with x1cry filters 
Message ID: 2137     Entry time: Fri Aug 17 14:56:59 2018     Reply to this: 2140
Author: aaron 
Type: Update 
Category: Cryo Q 
Subject: Finding Modes 

[Aaron, Mandy]

Yesterday at the cryo meeting, Mandy reported havnig trouble finding the eigemodes of the disk in air. Last night Aaron spent some time trying to find the modes, and left the diaggui running with a find swept sine TF measurement overnight; despite these efforts and the observation that the ESD could generate some excess power over noise when driving at a single frequency, the TF measurement did not reveal any modes (screenshot attached) and Aaron was unable to find them manually in the lab.

Today, we are putting the GeNS system in vacuum to try to reduce the amount of acoustic noise hiding our modes. We took out the two dewars from Hi'iaka and replaced the top with the jacket bottom from Anu; thus we cannot cool down, but hopefully have a tight vacuum.

We put the GeNS into jacket, but the initial pumpdown stopped at 100 torr or so, indicating a large leak. We were really fast and loose with the o-rings, so it's possible we just need to wipe them down, confirm that all ports actually have o-rings, and regrease the viton.

Despite this, it's possible that just having the system in a closed volume with less airflow will be sufficient to identify some modes, so we put a pause on vacuum efforts and Mandy is now aligning the beam into the cryo.

--

We realized we'd forgotten one of the o-rings; on replacing this o-ring, the vacuum gets to 5mtorr, which is still way too high, but makes some sense given the state of the o-rings and the in-vacuum components that have been sitting in-air for quite some time.

We continued aligning the beam with much difficulty, eventually resorting to moving the periscope closer to the cryostat so we could use a wider range of ingoing angles.

Attachment 1: InAirGeNSTF.png  58 kB  Uploaded Fri Aug 17 16:07:52 2018  | Hide | Hide all
InAirGeNSTF.png
ELOG V3.1.3-